OpenStudy (anonymous):

If the dimensions of a rectangle are cut in half, then the new area is one fourth of the original area? (not sure) o.o

OpenStudy (anonymous):

true eg given a rectangle 6 * 2 = area of 12 each side cut in half = 3 * 1 = area of 3